OpenStudy (anonymous):

Can someone help me break down the following expression: x=3 sec (theta) reduces the expression square root of x squared - 9 enclosed in paranthesis to the following answers: a. 3 sec (theta). b. 9 sec squared (theta). c. 9 tan squared (theta). d. 3 tan (theta).

OpenStudy (anonymous):

Am I in the right topic?

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

Howdy stranger ^.^ How can I be of service?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

do you understand how I wrote the problem?

OpenStudy (anonymous):

No. But for some reason, I think I already know the answer. Weird, huh? :D

OpenStudy (anonymous):

want me to write out the problem more clearly?

OpenStudy (anonymous):

Sure, why not?

ganeshie8 (ganeshie8):

\(\large \sqrt{x^2-9}\) \(\large x = 3 \sec \theta\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

simply substitute x value and simplify ?

OpenStudy (anonymous):

wat does 3 sec theta mean though?

OpenStudy (anonymous):

what ganeshie said. it all boils down to that :D

OpenStudy (anonymous):

but wait

OpenStudy (anonymous):

wat does 3 sec theta equal?

OpenStudy (anonymous):

x. LOL

ganeshie8 (ganeshie8):

it equals x

ganeshie8 (ganeshie8):

\(\large \sqrt{ (3 \sec \theta )^2-9}\)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[3\sec\theta= \frac3{\cos \theta }\]

OpenStudy (anonymous):

give me a moment

OpenStudy (anonymous):

3 sec theta minus 3?

ganeshie8 (ganeshie8):

\(\large \sqrt{ (3 \sec \theta )^2-9}\) \(\large \sqrt{9 \sec^2 \theta -9}\) \(\large \sqrt{9 (\sec^2 \theta -1)}\)

OpenStudy (anonymous):

don't forget the "squared" part

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

then? 9 is a perfect square, so it comes out of the radical (its square root anyway)

OpenStudy (anonymous):

so wait, would it be 3(sec theta+1)?

OpenStudy (anonymous):

don't forget the squared on the sec...and the square root

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

you think you could write it out for me?

OpenStudy (anonymous):

I'm still kinda confused at this part

OpenStudy (anonymous):

lol of course ^.^ \[\Large \sqrt{9(\sec^2\theta \ - 1)}= \sqrt9\sqrt{\sec^2\theta \ - 1}= 3\sqrt{\sec^2\theta \ - 1}\]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so the 9 comes out of the radical and makes itself a radical

OpenStudy (anonymous):

this is algebra stuff though :/

OpenStudy (anonymous):

I'm sorry bro

OpenStudy (anonymous):

I've never seen it done that way

OpenStudy (anonymous):

well back to trigonometry...i'm sure you can find a suitable way to simplify \[\sec^2 \theta - 1\]right?

OpenStudy (anonymous):

ahh yeah give me a moment

OpenStudy (anonymous):

that equation sec squared theta-1=tan squared theta

OpenStudy (anonymous):

so you're not totally lost after all :D \[\Large 3\sqrt{\sec^2\theta \ - 1}=3\sqrt{\tan^2\theta}\] Surely you can apply the finishing touch from here? ^.^

OpenStudy (anonymous):

and that would simply to 3 tan (theta)

OpenStudy (anonymous):

that does the trick, doesn't it? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I appreciate it

OpenStudy (anonymous):

do you think you could help me with another problem?

OpenStudy (anonymous):

depends on the problem :D